# ResAD: Normalized Residual Trajectory Modeling for End-to-End Autonomous Driving

### v1

@@ 3.3 Multimodal Trajectory Ranker
Inspired by VADv2 [3] and Hydra-MDP [23], we develop a Trajectory Ranker to select the optimal trajectory from multiple modalities by using the output from the planning model. Given a set of trajectory candidates $v_{k}$, where k is the vocabulary size, we feed them into a Transformer to facilitate interaction with the perception representations, $E_{env}$, which can be expressed as follows:
@@ 4.1 Benchmark
We evaluate the proposed ResAD on the NAVSIM v1 [8] and NAVSIM v2 [1] benchmark. NAVSIM is built upon the real-world NuPlan dataset [19] and exclusively features relevant annotations and sensor data sampled at 2 Hz. The NAVSIM dataset contains two parts: navtrain and navtest, including 1192 and 136 scenarios respectively, used for trainval and test.
@@
NAVSIM v2. In NAVSIM v2, a new Extended PDM Score (EPDMS) is introduced in NAVSIM v2 , which can be formulated as:
@@ 4.2 Implementation Details
For fair comparison, our model adopts an identical perception module and ResNet-34 backbone as Transfuser [7]. The model takes two types of input: three forward-facing camera images, which are individually cropped, downscaled, and then concatenated into a single 1024×256 tensor; and a rasterized BEV representation of the LiDAR point cloud.  ResAD is equipped with 2 cascaded diffusion layers. We set the mode number $K_{\mathrm{train}}=20$ for training and $K_{\mathrm{infer}}=200$ for testing. The model is trained from scratch on the navtrain split for 100 epochs using the DDPM, with a timestep $T$ of 1000. The training is distributed across 8 NVIDIA L20 GPUs, with a total batch size of 512, and is optimized using AdamW. The ranker’s training leverages a fixed trajectory vocabulary and the output from our frozen, pre-trained diffusion model to learn a scoring function. In inference, we use DDIM to sample the predictions with only 2 denoising steps. The resulting candidates are then evaluated by the trained ranker, which selects the highest-scoring trajectory as the output. We predict $T_{f}=8$ timesteps and the interval between each time step is 0.5s. For more details, please refer to the supplementary material.
@@ 4.3 Main Results
Quantitative Comparison. The results presented in Tab. 1 show that ResAD achieves a state-of-the-art performance on NAVSIM v1 navtest split, with a PDMS of 88.6. Our NC of 98.0 is on par with the highest scores, ensuring a high level of safety by minimizing collisions. The EP of 82.5 achieved by our model is a notable result, indicating efficient route completion. ResAD excels in DAC with a score of 97.3, outperforming WoTE’s 96.8. This suggests our model has a stronger adherence to lane boundaries and drivable areas, a critical aspect of safe and predictable driving behavior. On the more challenging NAVSIM v2 benchmark, the advantages of ResAD are further extended. As shown in Tab. 2,  ResAD achieves the best or second-best performance across almost all extended sub-metrics. Specifically,  ResAD achieves an EPDMS of 85.5, surpassing DiffusionDrive by 1.0. It achieves a higher EP score of 88.2 (vs. 87.5), indicating it completes routes more effectively. Furthermore, it shows a significant advantage in DAC with a score of 97.2 versus 95.9, confirming its ability to generate more precise trajectories that better adhere to lane boundaries.  ResAD also exhibits finer vehicle handling, with slightly better scores in LK.
@@
$\mathcal{M}_{4}$ $\mathcal{M}_{3}$ + IRP 98.0 97.3 82.5 94.2 100 88.6
Figure 3: Visual comparison of ResAD. DiffusionDrive relies on a static, context-agnostic vocabulary, often proposing infeasible trajectories (circled in red). In contrast, the proposed ResAD dynamically generates context-aware trajectories by perturbing the ego-vehicle’s velocity, addressing limitations in static approaches.
Qualitative Comparison. A qualitative comparison on NAVSIM (Fig. 3) highlights the different multimodal strategies of ResAD and DiffusionDrive. While both successfully avoid the mode collapse typical of vanilla diffusion, their underlying approaches diverge significantly. DiffusionDrive relies on a static, predefined trajectory vocabulary. This context-agnostic approach forces it to generate many irrelevant or unfeasible options, such as proceeding straight in a sharp turn scenario (highlighted by red circles in the figure). Although a subsequent filtering step can prune these invalid paths, this two-stage process is inherently inefficient. In contrast, ResAD overcomes this limitation through the distinct trajectory modeling shift. This is achieved through a mechanism of perturbing the ego-vehicle’s velocity. It directly explores a set of plausible behaviors, generating trajectories that are inherently consistent with the immediate driving context. A real-world vehicle demonstration of the proposed method is available in the supplementary material.
Figure 4: The impact of PRNorm on training efficiency and performance. This figure shows the loss and mean PDMS curves for  ResAD trained with or without the proposed PRNorm.
4.4 Ablation Studies
Component Analysis. To validate the efficacy of each proposed component in  ResAD, we conduct a comprehensive ablation study, with the results detailed in Table 3. Our analysis begins by integrating the Multimodal Trajectory Ranker module into the base model, forming $\mathcal{M}_{1}$. It offers only a slight performance benefit, suggesting poor multimodal ability as its outputs were mostly limited to a small area. The introduction of Trajectory Residual Modeling (TRM) significantly boosted performance, with the DAC metric improving from 94.3 to 96.6 and EP from 77.8 to 80.3, underscoring its role in improving path completion and safety. In addition, integrating PRNorm enhanced performance, particularly for EP, which demonstrates its value in normalizing feature representations and accelerating training. Finally, incorporating Inertia Reference Perturbation (IRP) to enhance multi-modal planning brought substantial gains, increasing the PDMS score from 87.2 to 88.6. The remarkable effectiveness of IRP is enabled by our Normalized Residual Trajectory Modeling approach. By deconstructing the trajectory data, our model can cleverly foster multi-modality without relying on a fixed trajectory vocabulary, allowing ResAD to generate a diverse set of trajectories that are better aligned with the current environmental state.
@@ 4.4 Ablation Studies
As shown in Fig. 4, PRNorm enables a significantly faster decline in loss compared to the baseline (vanilla min-max normalization), accelerating model convergence. Furthermore, we calculate the PDMS of the predicted trajectory every step, which is also higher with PRNorm. It demonstrates its comprehensive benefits to both training efficiency and the final performance.
@@
Effect of the Normalized Residual Trajectory Modeling. To further validate the effectiveness of our proposed Normalized Residual Trajectory Modeling, we conducted extensive experiments on two heterogeneous planning models. The results are shown in Tab. 4. Specifically, we evaluated it on Transfuser, which represents the MLP-based planning network, and TransfuserDP, which is an extension of Transfuser incorporating a UNet diffusion decoder, representing the diffusion-based planning network. Our findings consistently demonstrate that the proposed Normalized Residual Trajectory Modeling significantly enhances trajectory quality across both types of planning networks. The integration of the proposed TRM and PRNorm yields notable improvements across several crucial performance metrics. For the Transfuser baseline, the PDMS is improved from 84.0 to 85.4 with the help of TRM. With the engagement of PRNorm, the PDMS is further increased to 85.9. Consistent improvements are also observed with TransfuserDP. Consistent gains on diverse metrics validate Normalized Residual Trajectory Modeling as a generalizable and effective method for improving the safety and reliability of E2EAD systems.
@@ 5 Conclusion
In this work, we revisit the conventional future-trajectory-prediction paradigm in E2EAD. We argue that directly predicting a vehicle’s trajectory from sensor data forces models into a state of causal confusion and creates a planning horizon dilemma, undermining safety and reliability. Our proposed framework, ResAD, confronts these challenges by reframing the learning objective. By first establishing a deterministic inertial reference, we provide a strong physical prior that anchors the prediction task. The model then learns to predict the residual, i.e., the necessary deviations from this baseline, which encourages it to focus on the external causal factors, such as obstacles and traffic rules, that govern safe navigation. Furthermore, we introduced Point-wise Residual Normalization (PRNorm) to specifically tackle the optimization imbalance. PRNorm re-weights the learning objective at each waypoint, preventing large-magnitude errors from distant, uncertain predictions from dominating the training process and ensuring that critical, near-term adjustments are properly prioritized. Our state-of-the-art results on NAVSIM demonstrate that this conceptual shift significantly simplifies the learning task and provides a more robust foundation for future E2EAD systems.

### v2

@@ 3.3 Multimodal Trajectory Ranker
Inspired by VADv2 [chen2024vadv2] and Hydra-MDP [li2024hydra], we develop a Trajectory Ranker to select the optimal trajectory from multiple candidates generated by the planning module. Given a set of trajectory candidates $v_{k}$, where k is the number of trajectories, we feed them into a Transformer to facilitate interaction with the perception representations, $E_{env}$, which can be expressed as follows:
@@ 4.1 Benchmark
We evaluate the proposed ResAD on the NAVSIM v1 [dauner2024navsim] and NAVSIM v2 [Cao2025ARXIV] benchmark. NAVSIM is built upon the real-world NuPlan dataset [karnchanachari2024towards] and exclusively features relevant annotations and sensor data sampled at 2 Hz. The NAVSIM dataset contains two parts: navtrain and navtest, including 1192 and 136 scenarios respectively, used for trainval and test.
@@
NAVSIM v2. A new Extended PDM Score (EPDMS) is introduced in NAVSIM v2, which can be formulated as:
@@ 4.2 Implementation Details
For fair comparison, our model adopts an identical perception module and ResNet-34 [resnet] or V2-99 [v299] as its backbone following existing methods [chitta2022transfuser, li2025hydra++]. The proposed ResAD is equipped with 2 cascaded diffusion decoder layers. We set the number of modes $K_{\mathrm{train}}=20$ for training and $K_{\mathrm{infer}}=200$ for testing. The model is trained from scratch on the navtrain split for 100 epochs using the DDPM [ho2020denoising], with a timestep $T$ of 1000. The training is distributed across 8 NVIDIA L20 GPUs, with a total batch size of 512, and is optimized using AdamW. In inference, we use DDIM [songdenoising] to sample the predictions with only 2 denoising steps. The resulting candidates are then evaluated by the trained ranker, which selects the highest-scoring trajectory as the output. We predict $T_{f}=8$ timesteps and the interval between each time step is 0.5s. For more details, please refer to the supplementary material.
Figure 3: Visual comparison of ResAD. This figure compares the 20 trajectory candidates from ResAD and DiffusionDrive. DiffusionDrive relies on a static, context-agnostic vocabulary, often proposing infeasible or irrelevant trajectories (highlighted by red circles). In contrast, ResAD dynamically generates a set of context-aware trajectories via IRP. This demonstrates our method’s more efficient multimodal exploration, which avoids wasting capacity on invalid options and leads to a higher multimodal planning quality ($\mathcal{P}_{m}$), as validated in Sec. $4.$
@@ 4.3 Main Results
Quantitative Comparison. The results presented in Tab. 1 show that ResAD achieves a SOTA performance on the NAVSIM v1 navtest split, with a PDMS of 88.8 using a ResNet-34 backbone, surpassing strong baselines such as DiffusionDrive (PDMS 88.1) and WoTE (PDMS 88.3) under the same backbone. With a stronger V2-99 backbone, ResAD further improves to PDMS 90.6, surpassing GoalFlow and Hydra-MDP. On the more challenging NAVSIM v2 benchmark, the advantages of ResAD are further extended. As shown in Tab. 2ResAD reaches the best or second-best performance across almost all extended sub-metrics. Relative to DiffusionDrive,  ResAD improves notably on EP (88.2 vs. 87.5) and DAC (97.2 vs. 95.9), indicating stronger route completion and adherence to drivable areas under the richer EPDMS protocol.
Qualitative Comparison. A qualitative comparison on NAVSIM (Fig. 3) highlights the different multimodal strategies of ResAD and DiffusionDrive. While both successfully avoid the mode collapse typical of vanilla diffusion, their underlying approaches diverge significantly. DiffusionDrive relies on a static, predefined trajectory vocabulary. This context-agnostic approach forces it to generate many irrelevant or unfeasible options, such as proceeding straight in a sharp turn scenario (highlighted by red circles in the figure). Although a subsequent filtering step can prune these invalid paths, this two-stage process is inherently inefficient. In contrast, ResAD overcomes this limitation through its distinct trajectory modeling shift. This is achieved through a mechanism of perturbing the ego-vehicle’s velocity. It directly explores a set of plausible behaviors, generating trajectories that are inherently consistent with the immediate driving context. A real-world vehicle demonstration of the proposed method is available in the supplementary material.
4.4 Ablation Studies and Analysis
@@
$\mathcal{M}_{4}$ $\mathcal{M}_{3}$ + IRP 98.0 97.5 83.4 94.1 100 88.8
Figure 4: PRNorm stabilizes optimization and accelerates convergence. This figure compares the training dynamics of ResAD with and without PRNorm.
Component Analysis. As summarized in Tab. 3, adding the Ranker to the baseline ($\mathcal{M}_{0}\rightarrow\mathcal{M}_{1}$) yields only marginal benefit (PDMS +0.2) and mildly improves safety, indicating its choices concentrate in a narrow mode. Introducing Trajectory Residual Modeling (TRM) delivers the key boost (DAC +2.3, EP +2.5, PDMS +1.2). By providing a strong physical prior, TRM compels the model to learn the underlying driving logic rather than relying on the spurious correlations it might learn from raw trajectory data. This is further qualitatively validated in Fig. 5. Adding Point-wise Residual Normalization (PRNorm) further refines performance and stabilizes convergence, improving EP to 81.4 and PDMS to 87.2. PRNorm stabilizes optimization by resolving spatio-temporal scale variance, preventing far-field errors from dominating. This refines the trajectory quality, leading to another notable increase in EP (+1.1). Finally, Inertial Reference Perturbation (IRP) delivers the largest single improvement, reaching 88.8 PDMS (+1.6 over $\mathcal{M}_{3}$). IRP unlocks true, context-aware multimodality. It generates a diverse yet plausible set of intent hypotheses, providing high-quality candidates for the ranker to evaluate. This is the key that enables the model to excel in complex scenarios, dramatically improving EP (+2.0) and DAC (+0.8).
@@
Impact of PRNorm on Training Dynamics. As shown in Fig. 4, PRNorm enables a significantly faster decline in loss compared to the baseline (vanilla min-max normalization), accelerating model convergence. Furthermore, we calculate the PDMS of the predicted trajectory every step, which is also higher with PRNorm. It demonstrates its comprehensive benefits to both training efficiency and the final performance.
Qualitative Analysis of the TRM. As shown in Fig. 5, we remodel the trajectory as a residual against the inertial reference (IR). The IR can serve as a counterfactual baseline: in steady cruising, predictions largely coincide with the IR, yielding small residuals and stable optimization; in dynamic scenes (merging, intersections, sharp turns, braking), the baseline becomes highly inaccurate, producing strong, directional residuals. These residual vectors point away from obstacles or follow road curvature, indicating the model learns why to deviate (due to obstacles, yielding, or turning) rather than merely fitting a complex trajectory. Moreover, IRP samples a bundle of lightly perturbed baselines around the original IR, expanding the physically plausible neighborhood. The model learns sharper residuals over these references, avoiding mode collapse. Together, IR + IRP focus capacity on what must change and encourage exploration of a more effective trajectory space, enabling robust predictions with few denoising steps.
Figure 5: The impact of Inertial Reference. We highlight the 1st-ranked and 5th-ranked trajectory with their corresponding IRs. Red arrows are used to highlight key examples of the predicted residual.
@@
Table 5: Comparison of running time and performance. FPS and total runtime are measured on an NVIDIA 4090 GPU.
[…]
TransfuserDP 84.6 20 84.4 130.0ms 101M 7
DiffusionDrive 88.1 20 60.3 7.6ms 60M 45
ResAD w/o Ranker - 20 86.1 7.7ms 62M 45
ResAD 88.3 20 86.1 11.4ms 68M 37
Superiority of the Normalized Residual Trajectory Modeling. We demonstrate the numerous advantages of Normalized Residual Trajectory Modeling (NRTM) through further experiments: (1) Plug-and-Play. NRTM is a drop-in strategy that requires limited changes to the architecture or loss design: attach TRM (+PRNorm) to any planner and train as usual. As shown in Tab. 4, it consistently lifts performance on both an MLP planner (Transfuser) and a diffusion planner (TransfuserDP). On Transfuser, TRM raises PDMS from 84.0 to 85.2, and TRM+PRNorm further to 85.6; comparable gains are observed on TransfuserDP. By residual modeling, NRTM captures essential high-order driving primitives and, via spatio-temporal data alignment, accelerates optimization. It delivers method-agnostic improvements to E2EAD safety and efficiency with minimal integration cost. (2) Efficient and Context-aware Multimodal Planning. As shown in Tab. 5, compared with DiffusionDrive, all increases in parameters and inference time for ResAD come from the Ranker, which selects the optimal trajectory from a set of candidates. Beyond reporting PDMS, we also introduce the mean PDMS $\mathcal{P}_{m}$ to evaluate the average quality of the model’s multimodal trajectory outputs. ResAD attains a substantially higher $\mathcal{P}_{m}$, providing further quantitative evidence that ResAD generates context-aware multimodal trajectories. (3) Test-Time Scaling.
Figure 6: Dynamic number of ${K}_{\text{infer}}$. This figure plots PDMS as a function of the inference-time sampled noise numbers ${K}_{\text{infer}}$.
As shown in Fig. 6, while DiffusionDrive exhibits a distinct performance drop when $K_{\text{infer}}<K_{\text{train}}$ and remains around 88.1 when $K_{\text{infer}}>20$,  ResAD realizes a consistent performance and continuous growth. The introduction of IR provides effective physical prior, constraining predictions within a context-relevant region and effectively avoiding invalid samples. Therefore, NRTM not only boosts the performance but also converts extra inference compute into consistent PDMS gains, revealing a compute-adaptive, anytime advantage at deployment.
@@ 5 Conclusion
In this work, we revisit the conventional E2EAD paradigm, highlighting a key challenge where directly predicting raw trajectories can create a severe optimization imbalance and risks learning spurious correlations. We propose ResAD, a novel framework that fundamentally reframes the learning problem. Our approach introduces two key innovations: Trajectory Residual Modeling (TRM) and Point-wise Residual Normalization (PRNorm). TRM provides a strong physical prior via an inertial reference, forcing the model to learn only the necessary, context-driven residuals. This simplifies the task and focuses capacity on learning robust driving logic, not spurious patterns. PRNorm mitigates the optimization imbalance, ensuring that safety-critical, near-term adjustments are prioritized over uncertain long-term predictions. Our state-of-the-art results on NAVSIM are a direct consequence of this conceptual shift. By taming the imbalanced data and simplifying the learning objective, ResAD provides a significantly more robust, stable, and scalable foundation for E2EAD systems.